OpenStudy (anonymous):

solve for x: (x+4)/2=7

OpenStudy (ganpat):

x / (x +4) /2 = 7 2x / (x +4) = 7 2x = 7x + 28 -5x = 28 x = -28/5 ???

OpenStudy (waleed_imtiaz):

x+4=14. so, x=14-4=10

OpenStudy (anonymous):

thank you